(3 points) The Highway Safety Department wants to study the driving habits of individuals. A sample of 49 cars traveling on a particular stretch of highway revealed an average speed of 68.7 miles per hour with a standard deviation of 9.8 miles per hour. Round to 4 decimal places.

1.Calculate a 99% confidence interval for the true mean speed of all cars on this particular stretch of highway.

( , )

2. What sample size is needed to estimate the true average speed to within 0.5 mph at 99% confidence? Note: For consistency's sake, round your t* value to 3 decimal places before calculating the necessary sample size.

Choose n =
Answer #1

Solution :

Given that,

1) Point estimate = sample mean = \bar x = 68.7

sample standard deviation = s = 9.8

sample size = n = 49

Degrees of freedom = df = n - 1 = 49 - 1 = 48

At 99% confidence level

\alpha = 1 - 99%

\alpha =1 - 0.99 =0.01

\alpha/2 = 0.005

t\alpha/2,df = t0.005,48 = 2.682

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.682 * (9.8 / \sqrt 49)

Margin of error = E =3.75

The 99% confidence interval estimate of the population mean is,

\bar x  ± E  

= 68.7  ± 3.75

= ( 64.95, 72.45)

2) margin of error = E = 0.5

sample size = n = [t\alpha/2,df* s / E] 2

n = [2.682 * 9.8 / 0.5]2

n = 2763.31

Sample size = n = 2764